Listed below are the amounts of mercury​ (in parts per​ million, or​ ppm) found in tuna sushi sampled at different stores. The sample mean is 0.921 ppm and the sample standard deviation is 0.195 ppm. Use technology to construct a 90​% confidence interval estimate of the mean amount of mercury in the population

0.71

1.11

1.25

0.93

0.82

0.76

0.87

What is the confidence interval estimate of the mean amount of mercury in the​ population?

ppm<μ<ppm

​(Round to three decimal places as​ needed.)
Answer #1

Solution :

t /2,df = 1.895

Margin of error = E = t/2,df * (s /n)

= 1.895 * (0.195 / 7)

Margin of error = E = 0.140

The 90% confidence interval estimate of the population mean is,

- E < < + E

0.921 - 0.140 < < 0.921 + 0.140

0.781 ppm < < 1.061 ppm